# LLMs for Targeted Sentiment in News Headlines: Exploring Different Levels of Prompt Prescriptiveness

### v1

@@ 5 Conclusion
Building on parallels with annotation paradigms for subjective tasks, we investigated the performance of LLM in-context learning for targeted sentiment analysis on news headlines. Our findings indicate that, apart from few-shot prompting, predictive accuracy rises with prompt prescriptiveness level, though the optimal level varies by model. LLM uncertainty, while not correlating with human subjectivity, tends to be well-calibrated and also increases with prompt prescriptiveness level.
@@
We find several limitations in this work. Firstly, our choice of LLMs is restricted. This is primarily due to computing and budget constraints. We are aware that a more expansive collection of models is necessary for a more comprehensive overview of LLM performance, along with models larger than 7B parameters. Additionally, we prompted both GPT models using batches of data, which impacted performance during initial tests, but did not warrant the high costs of repeating the prompt for each individual instance.
@@ A.1 Additional Info on Datasets
A.2 Additional Information on GPU Usage
@@
A.3 Additional Information on Used Toolkits
@@
A.4 Prescriptiveness Levels Used in Prompts
Level 1: Descriptive without definitions, exploring the fundamental concepts of sentiment and targeted sentiment.
@@ A.4 Prescriptiveness Levels Used in Prompts
System: You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines. The available sentiment classes are positive, neutral, and negative. For each given headline, identify the targeted sentiment class towards the entity.
User: Classify targeted sentiment towards entity {entity} in the following news headline: {headline}
Level 2: Includes a definition of targeted sentiment specifically within the framework of news headlines.
@@
System: You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines. Targeted sentiment involves understanding the author’s intention to evoke emotion towards a target entity, considering the deliberate choice in conveying news and recognizing that the same information can be presented in various ways, with the understanding that such intentional choices aid in detecting the targeted sentiment. The available sentiment classes are positive, neutral, and negative. For each given headline, identify the targeted sentiment class towards the entity.
User: Classify targeted sentiment towards entity {entity} in the following news headline: {headline}
Level 3: Provided with concise guidelines.
@@
System: You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines. Targeted sentiment is the emotional stance the author aims to convey specifically towards a mentioned entity. It involves interpreting the intention behind the author’s choice of language and tone when discussing the target entity. The sentiment is not only influenced by the conveyed information but also by the author’s subjective evaluation and emotional coloring of the entity. Actions associated with the entity play a role in determining the sentiment, with negative actions implying a negative quality and, consequently, a negative sentiment. Distinguishing between negative actions and negative occurrences is crucial, as negative occurrences towards the entity don’t color the entity. In headlines featuring a quote, the entity authoring the quote is attributed neutral sentiment as they are merely conveying an opinion. The overall goal of the author, whether it be praise or criticism, is considered in cases of headlines with a mix of positive and negative views. In summary, targeted sentiment is the nuanced emotional evaluation directed specifically at a particular entity within the context of news reporting. The available sentiment classes are positive, neutral, and negative. For each given headline, identify the targeted sentiment class towards the entity.
User: Classify targeted sentiment towards entity {entity} in the following news headline: {headline}
Level 4: Comprehensive instructions provided as guidelines, excluding examples.
@@
System: You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines.
Guidelines for Targeted Sentiment Annotation:
@@
User: Classify targeted sentiment towards entity {entity} in the following news headline: {headline}
Level 5: Complete instructions presented as guidelines, inclusive of examples and detailed explanations.
System: You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines.
Guidelines for Targeted Sentiment Annotation:
1. Detecting Sentiment through Author’s Intent and News Presentation:
Evaluate the intended sentiment towards an entity by analyzing the emotions the author aims to evoke and recognizing that news can be conveyed in multiple ways, with the chosen manner of conveyance serving a purpose and aiding in targeted sentiment detection.
Examples Illustrating Sentiment towards Entity Solin:
@@ 1. Detecting Sentiment through Author’s Intent and News Presentation:
2. Impact of Entity Actions:
Recognize that entity actions play a role in shaping sentiment, particularly with negative actions like murder and theft suggesting a negative quality. However, distinguish between negative actions where the entity is the perpetrator and negative occurrences where the entity is the recipient. Acknowledge that in the case of negative occurrences, the entity cannot be held responsible for the consequences but may be in a negative situation as a result, implying neutrality in the assessment.
Headlines with negative quality of entities linked to their actions:
@@ 3. Neutrality of Quoting Authors:
Define sentiment towards the entity by considering the author’s stance in a statement, whether the author is the headline writer or the individual quoted. When conveying someone’s sentiment in a quote, transfer that sentiment to the mentioned entity. In headlines quoting individuals, recognize two entity types: the statement’s author and the entities mentioned in the quote. If the target entities in the headline are the authors of the statement, the sentiment is typically neutral since they serve as conveyors of an opinion.
Examples of Handling Quotes in Headlines:
@@
4. Overall Authorial Goal:
Consider the author’s overall goal, whether it involves praise or criticism, especially in mixed-view headlines.
Example of a Combined Statement (Combination of Positive and Negative Views)
@@ 4. Overall Authorial Goal:
User: Classify targeted sentiment towards entity {entity} in the following news headline: {headline}
Level 6: Full set of annotator instructions
System: You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines. Here are some guidelines for detecting targeted sentiment in news headlines:
To determine sentiment towards an entity, we consider the kind of emotion the statement’s author intended to evoke regarding the target entity, that is, how the author intended to "color" that entity. To aid in discerning the intended sentiment towards the entity, one can consider the fact that the same piece of news can always be conveyed in multiple ways. The chosen manner of conveying a piece of news is selected with a purpose, and understanding that intention can be utilized for targeted sentiment detection.
An example of various ways of reporting the same news about entity Solin:
Targeted Sentiment: Negative
Explanation: Negative sentiment is attributed to Solin due to the author’s intention to criticize the institution for the shortage of pediatricians.
Targeted Sentiment: Negative
Explanation: Similar negative sentiment is conveyed towards Solin by criticizing the shortage of medical staff.
Headline: ’U Solinu nastupio nedostatak liječničkog kadra’
Targeted Sentiment: Neutral
Explanation: Neutral sentiment is assigned as the author’s intention is to convey information without criticizing the responsible institutions.
@@
When detecting targeted sentiment, we can assign a quality to the target entity as an aid in determining sentiment, based on the emotion the statement’s author associates with it. The quality of the entity is linked to the actions of that entity, which can be either negative or positive. Negative actions of the entity, such as murder, theft, and other illegal or socially unacceptable activities like insults, are attributed to the quality of that entity. Negative actions signify a negative quality of the entity, implying a negative sentiment. The same approach will be applied in cases of positive actions of the entity, indicating a positive sentiment towards the entity. It is necessary to distinguish between the negative actions of an entity and negative occurrences towards the entity. In the case of negative actions by the entity, the entity is the perpetrator and therefore responsible for that action. In the case of negative occurrences towards the entity, the entity is the recipient of the negative action and cannot be held responsible for the consequences of the action, although it may be in a negative situation as a result.
Examples of linking entity quality to actions:
Entity: Elizabeth Holmes
Targeted Sentiment: Negative
Entity: Đuro Brodarac
Targeted Sentiment: Negative
Examples of negative occurences towards the entity.
Entity: Indonezija
Targeted Sentiment: Neutral
Entity: Mogadišu
Targeted Sentiment: Neutral
@@
We define sentiment towards the entity as the author’s stance towards the target entity in a statement. The statement’s author can be the person who wrote the article headline or the author whose quote is conveyed in the form of the article headline. When conveying someone’s negative/positive sentiment in a quote or paraphrase, that sentiment is transferred to the entity. In headlines conveying someone’s quote, there are two types of entities - the statement’s author and the entities mentioned in the quote. If the target entities in the headline are the authors of the statement, the sentiment towards them will usually be neutral because, in this case, they are just conveyors of an opinion. An exception is the following example with entity Gotovac, where the statement’s author comments on himself, and the expressed sentiment is then transferred to the author himself.
Examples of Handling Quotes in Headlines:
Headline: ’Milanović: Žao mi je što sam podržao Bidena’
Entity: Milanović
Targeted Sentiment: Neutral
Entity: Biden
Targeted Sentiment: Negative
Entity: Gotovac
Targeted Sentiment: Positive
Entity: Grbin
Targeted Sentiment: Negative
Entity: Anka Mrak Taritaš
Targeted Sentiment: Neutral
Entity: Tomašević
Targeted Sentiment: Negative
In the case of a headline containing a combination of positive and negative views towards the entity, the final goal of the author towards the entity is considered, i.e., whether the author aimed for praise or criticism.
Example of a Combined Statement (Combination of Positive and Negative Views):
Entity: Vanna
Targeted Sentiment: Positive
Explanation: Positive sentiment is assigned to Vanna as the author’s intention is to praise the improvement in her style despite also criticizing her previous dressing choices.
@@
User: Classify targeted sentiment towards entity {entity} in the following news headline: {headline}
A.5 Prompts for Modeling LLM Uncertainty
Distribution sampling
@@ A.5 Prompts for Modeling LLM Uncertainty
User: Your task is to imagine you are representing 6 different people detecting the targeted sentiment in Croatian news headlines, each following the given guidelines. For a headline and an entity, you need to return detected targeted sentiment for each of the 6 voters.
Detect targeted sentiment for entity ’entity’ in headline: ’headline’. Possible sentiment classes are positive, neutral and negative. Please return the answer in JSON format like:
["targeted sentiment 1":"class 1"
"targeted sentiment 2":"class 2"
"targeted sentiment 3":"class 3"
"targeted sentiment 4":"class 4"
"targeted sentiment 5":"class 5"
"targeted sentiment 6":"class 6"]
Verbal confidence assessment
@@
User: You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines. Following the given guidelines, please return the confidence for detection of each class.
Detect targeted sentiment for entity {entity} in headline: {headline}. Possible sentiment classes are positive, neutral and negative.
Please return the confidence for each class in format like:
["confidence positive class", "confidence neutral class" ,"confidence negative class"]
A.6 Measurements

### v3

@@ 5 Conclusion
Building on parallels with annotation paradigms for subjective tasks, we investigated the performance of LLM in-context learning for targeted sentiment analysis on news headlines. Our findings indicate that predictive accuracy increases with prompt prescriptiveness, though the optimal level varies by model, and only some models benefit from few-shot prompting. Calibration generally improves with prompt prescriptiveness, and self-consistency sampling aligns best with human label variation.
@@
We find several limitations in this work. Firstly, our choice of LLMs is restricted. This is primarily due to computing and budget constraints. We are aware that a more expansive collection of models is necessary for a more comprehensive overview of LLM performance, along with open-source models larger than 8B parameters. Additionally, we prompted both GPT models using batches of data, which impacted performance during initial tests, but did not warrant the high costs of repeating the prompt for each individual instance.
@@ A.2 Additional Information on Models
In our experiments, we employed the following LLMs:
Neural Chat
 [^1] https://ollama.com/library/neural-chat (7B): A fine-tuned model based on Mistral [^2] https://mistral.ai/news/announcing-mistral-7b/ with good coverage of domain and language.
Llama 3 instruct (8B)
 [^3] https://ollama.com/library/llama3:instruct : Instruction-tuned models fine-tuned and optimized for dialogue/chat use cases that outperform many of the available open-source chat models on common benchmarks.
Phi-3 Mini instruct
 [^4] https://ollama.com/library/phi3:instruct/blobs/fa8235e5b48f (3.8B): Phi-3 Mini is a lightweight, state-of-the-art open model by Microsoft [^5] https://www.microsoft.com/, trained with a focus on high-quality and reasoning dense properties.
Gemma
666 https://ollama.com/library/gemma:instruct (8.5B): Gemma is a lightweight, state-of-the-art open model built by Google DeepMind. [^7] https://deepmind.google/
GPT-4 Turbo
 [^8] https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 (560B): Latest generation OpenAI [^9] https://openai.com/ model in time of running our experiments. We used the gpt-4-1106-preview model.
GPT-3.5
Turbo [^10] https://platform.openai.com/docs/models/gpt-3-5-turbo (175B): Released in 2023, faster and more affordable OpenAI model. We used the gpt-3.5-turbo-0125 model.
A.3 Additional Information on GPU Usage
@@
A.4 Additional Information on Used Toolkits
@@
A.5 Complete results
In this section, we present complete results for the prescriptiveness prompting. In Table 6 F1 scores are provided for all models and levels, and in Table 7 ECE is given per level and model. Figure 3 shows it graphically.
Level NC GPT 3.5 GPT 4 Llama 3 Gemma Phi-3
SCS DP VCA SCS DP VCA SCS DP VCA SCS DP VCA SCS DP VCA SCS DP VCA
1 60.3 48.4 52.4 60.9 61.5 47.1 66.0 64.9 59.5 53.4 46.4 34.7 48.4 40.0 42.7 43.6 38.4 48.6
2 62.0 50.1 54.9 61.4 62.9 46.9 64.4 64.9 61.6 50.1 46.9 30.9 48.8 38.8 42.1 40.6 34.6 48.1
3 63.5 50.5 53.9 67.6 62.5 53.9 69.9 66.4 63.2 52.9 54.5 44.4 55.8 44.7 47.9 44.2 44.9 44.4
4 63.7 47.3 58.7 64.8 62.8 50.6 70.2 66.9 66.9 51.9 52.6 41.3 53.1 41.1 43.1 43.6 47.2 39.9
5 60.4 52.9 57.6 63.9 59.6 53.1 66.8 68.6 65.0 60.6 55.5 55.0 59.3 58.7 48.9 49.4 48.9 45.6
6 62.8 53.5 60.1 66.5 56.3 55.4 68.2 64.7 64.9 61.9 50.7 52.1 59.4 53.5 46.7 46.3 45.5 48.8
Table 6: F1 scores for levels of detail in prompt and uncertainty quantification metrics.
Level NC GPT 3.5 GPT 4 Llama 3 Gemma Phi-3
SCS DP VCA SCS DP VCA SCS DP VCA SCS DP VCA SCS DP VCA SCS DP VCA
1 13.1 6.3 15.9 11.3 6.0 16.6 11.1 8.5 10.1 15.3 15.9 32.5 17.2 20.5 31.8 17.2 18.5 27.9
2 12.1 5.7 15.6 10.9 5.4 16.9 11.6 9.1 10.1 16.4 15.0 32.9 17.0 21.8 31.9 17.0 19.8 26.8
3 11.1 6.1 15.3 9.2 10.9 15.1 10.1 9.5 10.5 15.1 15.1 31.8 14.8 20.7 31.1 14.8 15.4 24.3
4 10.6 8.9 14.2 5.8 7.0 15.4 10.4 9.7 9.4 15.7 13.9 31.6 15.6 20.1 31.5 15.6 13.1 27.2
5 11.3 7.2 13.6 9.7 7.2 16.6 10.8 9.5 11.1 13.2 13.7 30.3 13.6 11.0 27.5 13.6 9.6 25.3
6 9.4 5.1 12.6 10.5 9.8 14.2 10.5 10.6 10.7 12.2 14.8 31.6 13.4 15.7 27.7 13.4 12.3 26.5
Table 7: Expected calibration error (ECE) for levels of detail in prompt and uncertainty quantification metrics.
A.6 Measurements
A.7 Prescriptiveness Levels Used in Prompts
In this section, the complete prompts system and user prompts are given in tables. The yellow highlight shows an expansion in text and information compared to the previous level.
Level Prompt
@@ A.7 Prescriptiveness Levels Used in Prompts
1 You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines. The available sentiment classes are positive, neutral, and negative. For each given headline, identify the targeted sentiment class towards the entity.
@@
2 You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines. Targeted sentiment involves understanding the author’s intention to evoke emotion towards a target entity, considering the deliberate choice in conveying news and recognizing that the same information can be presented in various ways, with the understanding that such intentional choices aid in detecting the targeted sentiment. The available sentiment classes are positive, neutral, and negative. For each given headline, identify the targeted sentiment class towards the entity.
@@
3 You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines. Targeted sentiment is the emotional stance the author aims to convey specifically towards a mentioned entity. It involves interpreting the intention behind the author’s choice of language and tone when discussing the target entity. The sentiment is not only influenced by the conveyed information but also by the author’s subjective evaluation and emotional coloring of the entity. Actions associated with the entity play a role in determining the sentiment, with negative actions implying a negative quality and, consequently, a negative sentiment. Distinguishing between negative actions and negative occurrences is crucial, as negative occurrences towards the entity don’t color the entity. In headlines featuring a quote, the entity authoring the quote is attributed neutral sentiment as they are merely conveying an opinion. The overall goal of the author, whether it be praise or criticism, is considered in cases of headlines with a mix of positive and negative views. In summary, targeted sentiment is the nuanced emotional evaluation directed specifically at a particular entity within the context of news reporting. The available sentiment classes are positive, neutral, and negative. For each given headline, identify the targeted sentiment class towards the entity.
Table 8: System prompts used for inference on the SToNe dataset.
Level Prompt
@@
4 You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines. Guidelines for Targeted Sentiment Annotation:
@@
5 You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines. Guidelines for Targeted Sentiment Annotation: 1. Detecting Sentiment through Author’s Intent and News Presentation: Evaluate the intended sentiment towards an entity by analyzing the emotions the author aims to evoke and recognizing that news can be conveyed in multiple ways, with the chosen manner of conveyance serving a purpose and aiding in targeted sentiment detection. Examples Illustrating Sentiment towards Entity Solin:
@@
Table 9: System prompts used for inference on the SToNe dataset.
Level Prompt
5 2. Impact of Entity Actions: Recognize that entity actions play a role in shaping sentiment, particularly with negative actions like murder and theft suggesting a negative quality. However, distinguish between negative actions where the entity is the perpetrator and negative occurrences where the entity is the recipient. Acknowledge that in the case of negative occurrences, the entity cannot be held responsible for the consequences but may be in a negative situation as a result, implying neutrality in the assessment. Headlines with negative quality of entities linked to their actions:
@@ 3. Neutrality of Quoting Authors:
Define sentiment towards the entity by considering the author’s stance in a statement, whether the author is the headline writer or the individual quoted. When conveying someone’s sentiment in a quote, transfer that sentiment to the mentioned entity. In headlines quoting individuals, recognize two entity types: the statement’s author and the entities mentioned in the quote. If the target entities in the headline are the authors of the statement, the sentiment is typically neutral since they serve as conveyors of an opinion. Examples of Handling Quotes in Headlines:
@@
Table 10: System prompts used for inference on the SToNe dataset.
Level Prompt
5 4. Overall Authorial Goal: Consider the author’s overall goal, whether it involves praise or criticism, especially in mixed-view headlines. Example of a Combined Statement (Combination of Positive and Negative Views)
@@
6 You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines. Here are some guidelines for detecting targeted sentiment in news headlines: To determine sentiment towards an entity, we consider the kind of emotion the statement’s author intended to evoke regarding the target entity, that is, how the author intended to "color" that entity. To aid in discerning the intended sentiment towards the entity, one can consider the fact that the same piece of news can always be conveyed in multiple ways. The chosen manner of conveying a piece of news is selected with a purpose, and understanding that intention can be utilized for targeted sentiment detection. An example of various ways of reporting the same news about entity Solin: Headline: ’SRAMOTA USolinuse djeca nemaju gdje liječiti, roditelji očajni’ Targeted Sentiment: Negative Explanation: Negative sentiment is attributed to Solin due to the author’s intention to criticize the institution for the shortage of pediatricians. Headline: ’U Solinu radi samo jedna pedijatrica, roditelji traže hitno rješenje’ Targeted Sentiment: Negative Explanation: Similar negative sentiment is conveyed towards Solin by criticizing the shortage of medical staff. Headline: ’U Solinu nastupio nedostatak liječničkog kadra’ Targeted Sentiment: Neutral Explanation: Neutral sentiment is assigned as the author’s intention is to convey information without criticizing the responsible institutions. When detecting targeted sentiment, we can assign a quality to the target entity as an aid in determining sentiment, based on the emotion the statement’s author associates with it. The quality of the entity is linked to the actions of that entity, which can be either negative or positive. Negative actions of the entity, such as murder, theft, and other illegal or socially unacceptable activities like insults, are attributed to the quality of that entity. Negative actions signify a negative quality of the entity, implying a negative sentiment. The same approach will be applied in cases of positive actions of the entity, indicating a positive sentiment towards the entity. It is necessary to distinguish between the negative actions of an entity and negative occurrences towards the entity. In the case of negative actions by the entity, the entity is the perpetrator and therefore responsible for that action. In the case of negative occurrences towards the entity, the entity is the recipient of the negative action and cannot be held responsible for the consequences of the action, although it may be in a negative situation as a result. Examples of linking entity quality to actions: Headline: ’Bivša tehnološka direktorica Elizabeth Holmes osuđena na 11 godina zatvora’ Entity: Elizabeth Holmes Targeted Sentiment: Negative Explanation: Negative sentiment is assigned to Elizabeth Holmes based on her negative actions.
Table 11: System prompts used for inference on the SToNe dataset.
Level Prompt
@@
6 Headline: ’Zbog ubojstva srpskih civila sudit će se Đuri Brodarcu, bivšem Sanaderovom savjetniku’ Entity: Đuro Brodarac Targeted Sentiment: Negative Explanation: Negative sentiment is assigned to Đuro Brodarac due to his association with a serious crime. Examples of negative occurences towards the entity. Headline: ’Potres u Indoneziji: Poginulo najmanje 46 ljudi, ozlijeđenih oko 700’ Entity: Indonezija Targeted Sentiment: Neutral Explanation: Neutral sentiment is assigned to Indonesia as the entity is a recipient of a negative occurrence. Headline: ’Horor u Mogadišuu: U terorističkom napadu na hotel 10 mrtvih, ozlijeđen i somalijski ministar’ Entity: Mogadišu Targeted Sentiment: Neutral Explanation: Similar to the previous example, neutral sentiment is assigned to Mogadishu as it is a recipient of a negative occurrence. We define sentiment towards the entity as the author’s stance towards the target entity in a statement. The statement’s author can be the person who wrote the article headline or the author whose quote is conveyed in the form of the article headline. When conveying someone’s negative/positive sentiment in a quote or paraphrase, that sentiment is transferred to the entity. In headlines conveying someone’s quote, there are two types of entities - the statement’s author and the entities mentioned in the quote. If the target entities in the headline are the authors of the statement, the sentiment towards them will usually be neutral because, in this case, they are just conveyors of an opinion. An exception is the following example with entity Gotovac, where the statement’s author comments on himself, and the expressed sentiment is then transferred to the author himself. Examples of Handling Quotes in Headlines: Headline: ’Milanović: Žao mi je što sam podržao Bidena’ Entity: Milanović Targeted Sentiment: Neutral Entity: Biden Targeted Sentiment: Negative Explanation: Neutral sentiment is assigned to Milanović, who is conveying an opinion, while negative sentiment is assigned to Biden based on the conveyed sentiment. Headline: ’Gotovac: Ako sam ja politički antitalent, onda je tom antitalentu išlo bolje nego Grbinu’ Entity: Gotovac Targeted Sentiment: Positive Entity: Grbin Targeted Sentiment: Negative Explanation: Positive sentiment is assigned to Gotovac, who comments on himself, while negative sentiment is assigned to Grbin based on the conveyed sentiment. Headline: ’Anka Mrak Taritaš: Tužna sam i razočarana situacijom u Zagrebu. Tomašević ne bi dobio dobru ocjenu’ Entity: Anka Mrak Taritaš Targeted Sentiment: Neutral Entity: Tomašević Targeted Sentiment: Negative Explanation: Neutral sentiment is assigned to Anka Mrak Taritaš, the quoted individual, while negative sentiment is assigned to Tomašević based on the conveyed sentiment.
Table 12: System prompts used for inference on the SToNe dataset.
Level Prompt
6 In the case of a headline containing a combination of positive and negative views towards the entity, the final goal of the author towards the entity is considered, i.e., whether the author aimed for praise or criticism. Example of a Combined Statement (Combination of Positive and Negative Views): Headline: ’Vanna je definitivno promijenila stil naglavačke i dosadne kombinacije zamijenila onima koje prate trendove’ Entity: Vanna Targeted Sentiment: Positive Explanation: Positive sentiment is assigned to Vanna as the author’s intention is to praise the improvement in her style despite also criticizing her previous dressing choices. The available sentiment classes are positive, neutral, and negative. For each given headline, identify the targeted sentiment class towards the entity.
Table 13: System prompts used for inference on the SToNe dataset.
Prompt
@@
SCS Classify targeted sentiment towards entity {entity} in the following news headline: {headline}
@@
DP Your task is to imagine you are representing 6 different people detecting the targeted sentiment in Croatian news headlines, each following the given guidelines. For a headline and an entity, you need to return detected targeted sentiment for each of the 6 voters. Detect targeted sentiment for entity ’entity’ in headline: ’headline’. Possible sentiment classes are positive, neutral and negative. Please return the answer in JSON format like: ["targeted sentiment 1":"class 1" "targeted sentiment 2":"class 2" "targeted sentiment 3":"class 3" "targeted sentiment 4":"class 4" "targeted sentiment 5":"class 5" "targeted sentiment 6":"class 6"]
@@
VCA You are a helpful assistant who performs targeted sentiment classification in Croatian news headlines. Following the given guidelines, please return the confidence for detection of each class. Detect targeted sentiment for entity {entity} in headline: {headline}. Possible sentiment classes are positive, neutral and negative. Please return the confidence for each class in format like: ["confidence positive class", "confidence neutral class" ,"confidence negative class"]
Table 14: User prompt used for inference on the SToNe dataset accross methods for uncertainty quantification.